OpenStudy (anonymous):

find the slope of the following lines just by examination. 2x+3y=11 i dont get what they mean to do it by examination?

OpenStudy (anonymous):

will give medal!

OpenStudy (anonymous):

@iambatman do u know?

OpenStudy (aravindg):

2x+3y=11 2x+3y-11=0 For a line Ax+By+C=0 Slope=-A/B

OpenStudy (aravindg):

Can you get the answer now?